IN TRIANGLE ABC  2acsin1/2(A-B+C)=

Grade:11

1 Answers

Lab Bhattacharjee
8 years ago
A-B+C=\pi-2B\\ \implies\sin\dfrac{A-B+C}2=\cos B\\ \text{Now, }\cos B=\dfrac{c^2+a^2-b^2}{2ca}